Product of[0 -1] and [7 -2]
[-4 -2] [-1 0]

Answer by lainey(7)

Product of[0 -1] and [7 -2]
[-4 -2] [-1 0]
1) 0-1 = -1
2) 7-2 = 5
3) -4 - 2 = -6
4) -1 + 0 = -1
5) -1 + 5 = 4
6) 4 - 6 = -2
7) -2 -1 = -3

Solution is -3.